OpenStudy (anonymous):

solve \[\int\limits_{0}^{b} \sqrt{1-2\cos \theta}d \theta \]

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=integrate+sqrt%28+1+-+2cos+x%29 guess not bounded or what it is called

OpenStudy (anonymous):

but it's a definite integral....

OpenStudy (experimentx):

b does not have definite value .. also for 2cos\theta>1, it is not defined

OpenStudy (anonymous):

b=2Pi

OpenStudy (experimentx):

it's definitely going to have complex value

OpenStudy (anonymous):

maybe i made a mistake somewhere......

OpenStudy (anonymous):

thx anyway...

OpenStudy (experimentx):

yw

OpenStudy (anonymous):

ya i made a mistake, it should be: \[\int\limits_{0}^{2\pi} \sqrt{2-2\cos \theta}=\sqrt{2} \sqrt{2} \int\limits_{0}^{2\pi}\sin (\theta/2)d \theta \] :) @experimentX

OpenStudy (experimentx):

oh .. that simplifies it i guess the previous ones are called http://en.wikipedia.org/wiki/Nonelementary_integral

OpenStudy (anonymous):

ya, maybe.

OpenStudy (anonymous):

eliptical integrals might solve that, but that's not for me...:)

OpenStudy (experimentx):

well ... i don't have much info on it either :)